OpenStudy (anonymous):

solve the following using la place d^2x/dt^2 +2dx/dt +5x = 1 given t =0 , x =0 and dx/dt =0

OpenStudy (accessdenied):

Have you done any work with this one yet?

OpenStudy (anonymous):

yeah i have but the problem i dont know how to apply the 3rd condition dx/dt =0

OpenStudy (accessdenied):

dx/dt (0) = 0 becomes useful when you use properties of Laplace transform to convert the transform of the second derivative of x.

OpenStudy (accessdenied):

\( \displaystyle \mathcal{L} \left[ \frac{d^2 x}{dt^2 } \right] = s \mathcal{L} \left[ \frac{dx}{dt} \right] - \frac{dx}{dt} (0) \)

OpenStudy (accessdenied):

This just being a consequence of the case with a first derivative.

OpenStudy (anonymous):

so can u show me how in my equation

OpenStudy (accessdenied):

Well, you take the Laplace transform of both sides to start. \( \displaystyle \mathcal{L} \left[ \frac{d^2 x}{dt^2} + 2 \frac{dx}{dt} + 5x \right] = \mathcal{L} \left[ 1 \right] \) Linearity of the Laplace transform operator lets you take the Laplace transform term-by-term and factor out constants freely. \( \displaystyle \mathcal{L} \left[ \frac{d^2 x}{dt ^2 } \right] + 2 \mathcal{L} \left[ \frac{dx}{dt} \right] + 5 \mathcal{L} \left[ x \right] = \mathcal{L} \left[1 \right] \) That part is clear, correct?

OpenStudy (anonymous):

yeah

OpenStudy (accessdenied):

This property is then useful to us, concerning the derivatives of x: \( \displaystyle \mathcal{L} \left[ \frac{dx}{dt} \right] = s \; \mathcal{L} \left[ x \right] - x(0) \)

OpenStudy (accessdenied):

We can use it for the second derivative as well, simply calling the second derivative "the derivative of dx/dt". \( \displaystyle \mathcal{L} \left[ \frac{d^2 x }{dt^2 } \right] = s \mathcal{L} \left[ \frac{dx}{dt} \right] - \frac{dx}{dt} (0) \) and then using the original property for the laplace transform of dx/dt. \( \displaystyle = s \left( s \mathcal{L} \left[ x \right] - x(0) \right) - x'(0) \) \( \displaystyle = s^2 \mathcal{L} \left[ x \right] - s x(0) - x'(0) \) Does all of that seem clear then?

OpenStudy (anonymous):

yeah it is clear

OpenStudy (accessdenied):

Alright, so we can directly substitute those two cases into what we have above: \( \displaystyle \mathcal{L} \left[ \frac{d^2 x}{dt ^2} \right] + 2 \mathcal{L} \left[ \frac{dx}{dt} \right] + 5 \mathcal{L} \left[ x \right] = \mathcal{L}\left[ 1 \right] \) \( \displaystyle (s^2 \mathcal{L} \left[x \right] - s x(0) - x'(0) )+ 2 (s \mathcal{L} \left[ x \right] - x(0)) + 5 \mathcal{L} \left[ x \right] = \frac{1}{s} \) 1/s is the Laplace transform of 1. If this all makes sense, you should be able to substitute all the conditions and get the answer. :)

OpenStudy (anonymous):

all this it is make sense for me but how can apply the dx/dt = 0

OpenStudy (accessdenied):

x'(0) = 0 is the same as dx/dt = 0 at t=0. Just a bit of ease of notation.

OpenStudy (accessdenied):

When they say, "given t =0 , x =0 and dx/dt =0" x = 0 at t = 0 x(0) = 0 and dx/dt = 0 at t = 0. x ' (0) = 0.

OpenStudy (anonymous):

okayyyyyyyyyyyyyyyyyyyyy now i got it tanxxxxxxx so much sir

OpenStudy (accessdenied):

You're welcome! :)